HETAV PATEL

Grade 11,

What is the maximum and minimum possible resultant of two forces 2N and 1N?

What is the maximum and minimum possible resultant of two forces 2N and 1N?
 

Grade:11

1 Answers

Unicorn
2 years ago
Resultant of two qty A and B =root (A^2+B^2+2A.Bcos angle)
Maximum =root (4+1+2.2.1cos0)
                      =root9=3
Minimum when theta =90 degrees 
                     =root(4+1+2.2.1.0)
                     =root5